OpenStudy (ilovepurple42043):

i need help woth these sample questions. will you show me step-by-step how to do each of them? 1: x^2 + 15x + 44 2: n^2 - 15n + 56 3: x^2 - 4x - 12 4: the area of a rectangular rug is given by the trinomonial r^2 - 6r - 55.What are the possible demensions of the rug?What are the length and the idth of the rug? 5: x^2 + 2xy - 99y^2 6: 9y^2 + 27y + 14 7; 2x^2 + 3x - 9 8: the area of knitted blanket is 10x^2 - 11x - 6.What are the possible demensions of the blanket? 9: 25x^2 35x - 30 10: x2 - 22x + 12r 11: x^2 - 25 12: 64c^2 - 9 13: 144y^2 - 196 14: 9z^8 + 81z^2 - 7^2 - 63 15: v= 4

OpenStudy (ilovepurple42043):

v=4y^3 + 15y^2 + 9y

OpenStudy (tkhunny):

1: x^2 + 15x + 44 Is this an exercise in factoring? Fidn all the factorizations of 44 1 * 44 2 * 22 4 * 11 Everything is positive, so which of those factorizations, if any, when added together, gives 15? 1 + 44 = 45 Nope 2 + 22 = 24 Nope 4 + 11 = 15 -- We have a winner! (x+4)(x+11) Done.

OpenStudy (ilovepurple42043):

that was right and i understad a little more now thank you, can you help me with the others? @tkhunny

OpenStudy (tkhunny):

You can show me the next one. Go!

OpenStudy (ilovepurple42043):

n^2 - 15n + 56 @tkhunny

OpenStudy (anonymous):

Here is may be the fastest method (the Diagonal Sum Method) to solve these type of equations, when a = 1. First, recall the Rule of signs. When a and c have different signs, the roots have different signs When a and c have same sign, the roots have same sign - When a and b have different signs, both roots are positive - When a and b have same sign, both roots are negative. 1. Solve: x^2 + 15x + 44 = 0. Solving results in finding 2 numbers knowing their sum (-b) and their product (c). Both roots are negative (Rule of signs). Compose factor pairs of c = 44. Proceeding: (-1, -44)(-2, -22)(-4, -11). The last sum is -4 - 11 = -15 = -b. Then the 2 real roots are -4 and -11. Write in factoring form: (x + 4)(x + 11) = 0 2. Solve x^2 - 15x + 56 = 0. Both roots are positive. Compose factor pairs of c = 56. Proceeding: (1, 56)(2, 28)(7, 8). This last sum is 7 + 8 = 15 = -b. Then, the 2 real roots are 7 and 8. Factoring form: (x -7)(x - 8) = 0 3. Solve: x^2 - 4x - 12 = 0, Roots have different signs. Compose factor pairs of c = -12. Proceeding: (-1, 12)(-2, 6). This last sum is 4 = -b. Then the 2 real roots are: -2 and 6.

OpenStudy (anonymous):

5. x^2 + 2xy - 99y^2 . To factor, find 2 numbers b1 and b2 that satisfy: b1*b2 = -99y^2, and (b1 + b2) = 2. Compose factor pairs of -99y^2. Proceeding" ...(-9, + 11). They are b1 = -9 and b2 = 11. Replace in the equation the term 2yx by the 2 terms 11y and -9y then factor. Finally: (x - 9y)(x + 11y) 6. 9y^2 + 27y + 14.Proceed with the same method. b1*b2 = 126; (b1 + b2) = 27. b1 = 6; b2 = 21. (3y + 2)(3y + 7) 7. 2x^2 + 3x -9 = (2x - 3)(x + 3). b1*b2 = -18;(b1 + b2) = 3; b1 = -3, b2 = 6. 8. 10x^2 - 11x - 6 = (5x + 2)(2x - 3)

OpenStudy (ilovepurple42043):

@myininaya

OpenStudy (anonymous):

10x^2 - 11x - 6 = (5x + 2)(2x - 3). Possible dimensions of the blanket: x = 2 -> Dimensions: 12 x 1 x = 3 -> id 17 x 3 x = 4 -> id 22 x 5....

OpenStudy (anonymous):

4. Area of the rug: r^2 - 6r - 55 = (r + 5)(r - 11). Possible dimensions: r = 6 -> Dimension 1 x 12 r = r -> id 2 x 13 r = 8 -> id 3 x 14 ......